OpenStudy (anonymous):

solve x2+6x=-7 by completing the square!

OpenStudy (campbell_st):

ax^2 +bx + c to complete the square is (b/2) and add to both sides (6/2)^2 = 9 x^2 + 6x + 9 = -7 + 9 (x +3)^2 = 2 now take the square root of both sides \[x + 3 = \pm\sqrt{2}\] subtract 3 from both sides \[x = -3 \pm \sqrt{2}\]

OpenStudy (anonymous):

thanks!